Solve Problems Using Linear Systems
15. Some students at L.C.V.I. held a bake sale recently to raise money for a field trip. They charged $7 for fruit pies and $10 for meat pies. They sold a total of 52 pies and earned $424. How many of each type of pie did they sell?
(The answer is: 32 fruit pies, 20 meat pies, but i dont know how to solve it?)

Answer by checkley77(12844)

10M+7F=424
M+F=52 OR M=52-F
NOW REPLACE M WITH (52-F) IN THE FIRST EQUATION & SOLVE FOR F.
10(52-F)+7F=424
520-10F+7F=424
-3F=424-520
-3F=-96
F=-96/-3
F=32 FRUIT PIES WERE SOLD.
52-32=20 MEAT PIES WERE SOLD.
PROOF:
10*20+7*32=424
200+224=424
424=424